# ComputerRL: Scaling End-to-End Online Reinforcement Learning for Computer Use Agents

### v1

@@ Abstract
We introduce ComputerRL, a framework for autonomous desktop intelligence that enables agents to operate complex digital workspaces skillfully. ComputerRL features the API-GUI paradigm, which unifies programmatic API calls and direct GUI interaction to address the inherent mismatch between machine agents and human-centric desktop environments. Scaling end-to-end RL training is crucial for improvement and generalization across diverse desktop tasks, yet remains challenging due to environmental inefficiency and instability in extended training. To support scalable and robust training, we develop a distributed RL infrastructure capable of orchestrating thousands of parallel virtual desktop environments to accelerate large-scale online RL. Furthermore, we propose Entropulse, a training strategy that alternates reinforcement learning with supervised fine-tuning, effectively mitigating entropy collapse during extended training runs. We employ ComputerRL on open models GLM-4-9B-0414 and Qwen2.5-14B, and evaluate them on the OSWorld benchmark. The AutoGLM-OS-9B based on GLM-4-9B-0414 achieves a new state-of-the-art accuracy of 48.1%, demonstrating significant improvements for general agents in desktop automation. The algorithm and framework are adopted in building AutoGLM (Liu et al., 2024a).
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; GLM et al., 2024; Zeng et al., 2022; Team et al., 2023; DeepSeek-AI et al., 2025; Bai et al., 2023a; Yang et al., 2025) have dramatically expanded the scope and depth of artificial intelligence (AI) capabilities, driving a profound re-examination of our understanding of machine intelligence. Among all scenarios, the emergence of LLM-based GUI (graphical user interfaces) agents, capable of independently perceiving, reasoning, and executing complex tasks on user devices, has aroused particular interests from researchers (Xi et al., 2023; Wang et al., 2023; Liu et al., 2023). Given that desktops remain central to intelligence-intensive tasks, developing effective computer use agents is crucial for fundamentally transforming human-computer interactions (Agashe et al., 2025; Wu et al., 2024) and lifting capabilities of AI to higher levels.
@@
Despite previous attempts to develop computer use agents (Agashe et al., 2025; Lei et al., 2024; Xie et al., 2025), enabling them to operate autonomously over extended periods in real-world scenarios remains a significant challenge. The first primary obstacle arises from the fact that GUIs are inherently designed for human interaction, making the simulation of human actions by GUI agents (Liu et al., 2024b; OpenAI, 2025; Qin et al., 2025) a non-trivial and cumbersome endeavor. Second, current mainstream approaches of behavior cloning (BC) (Bain & Sammut, 1995; Bratko et al., 1995), including manual annotation (He et al., 2025) and model distillation (Sun et al., 2024; Xu et al., 2024), are limited in scalability and effectiveness. Manual annotation, while precise, is prohibitively labor-intensive for complex tasks. Model distillation, on the other hand, is constrained by the performance of the teacher models, limiting overall capability. Both methods typically exhibit poor generalization and limited error recovery abilities. Finally, although reinforcement learning (RL) has shown potential for desktop automation tasks (Lu et al., 2025; Feng et al., 2025), its practical application remains restricted due to computational complexity and methodological challenges. Complex environment, slow convergence, and known inefficiencies in RL training (Xie et al., 2024; Bonatti et al., 2024; Yu et al., 2025; Fu et al., 2025) severely limit its large-scale adoption in training computer use agents.
@@
In this work, we propose ComputerRL, an end-to-end algorithmic framework designed to advance desktop-level planning, reasoning, and device operation, including a new API-GUI interaction paradigm, a scalable RL training infrastructure for computer environment, and a RL algorithm for extended effective training. First, we introduce API-GUI, a large-scale, automatically constructed API ecosystem that enables the agent to transcend the inherent biases of human-oriented operational paradigms. It instead leverages a more machine-oriented approach for device interaction, which combines API calls and GUI actions and thereby significantly enhances both the versatility and overall performance of the agent. Second, we develop a distributed training infrastructure utilizing virtual machine clusters based on Docker and gRPC protocols for scalability, which is fully compatible with AgentBench (Liu et al., 2023). This infrastructure supports thousands of parallel environments, ensuring high scalability and consistent interactions across all environments. Additionally, we combine the training infrastructure with the AgentRL framework (Zhang et al., 2025) for efficient asynchronous training, substantially accelerating the training process. Finally, to counteract stagnation and convergence issues in extended RL training—specifically, entropy collapse and rising KL divergence—we propose Entropulse, which alternates between RL and periodic SFT phases. This approach maintains exploratory capacity and ensures continuous performance gains (Figure 1(b)).
@@
As a result, by harnessing end-to-end RL and optimization in the desktop environment, ComputerRL has achieved remarkable improvement in understanding and operating GUIs. Evaluation on the OSWorld benchmark (Xie et al., 2024) shows ComputerRL’s significant improvements (see Figure 1(a)) in computer use challenges, achieving a success rate of 48.1% (with 64% performance gain from RL), outperforming other state-of-the-art models including OpenAI CUA o3 (42.9%), UI-TARS-1.5 (42.5%), and Anthropic Claude Sonnet 4 (30.7%).
@@
  • We propose a new interaction paradigm, shift from human-centric to machine-oriented interaction frameworks by introducing a large-scale, automatically constructed API ecosystem integrated with conventional GUI operations. This approach addresses the inherent mismatch between human-designed interfaces and artificial agent capabilities, while achieving superior operational efficiency and generalization performance on computer-based tasks.
@@
  • We introduce Entropulse, a novel training methodology that systematically addresses the challenges of entropy collapse and KL divergence accumulation in extended RL training through strategic alternation between RL and SFT phases, enabling sustained performance improvements and achieving state-of-the-art performance on the OSWorld benchmark.
@@ 2 The ComputerRL Framework
The human-oriented design of GUI environments leads to inefficient interactions for agents, which impedes effective operation and learning. Furthermore, limited environment scalability poses a significant barrier to large-scale agent training. In this section, we present the design and implementation of the ComputerRL framework, as in Figure 3. We propose an API-GUI paradigm powered by LLMs, which integrates human-like GUI interactions with efficient API invocation. Furthermore, we develop a robust and highly scalable Ubuntu Desktop environment designed for large-scale parallel experimentation. To facilitate efficient model training, we use a fully asynchronous RL framework (Zhang et al., 2025).
Figure 3: Overview of ComputerRL framework. We introduce an API-GUI action paradigm that seamlessly integrates automatically constructed APIs with GUI actions to improve agent efficiency and effectiveness. A large-scale parallel desktop environment with over 1,000 real-world instances, combined with an asynchronous RL framework, enables efficient sampling and robust agent training.
@@ 2.1 General API-GUI Paradigm
A significant challenge confronted by existing GUI agents arises from their reliance on human-like GUI interactions for device manipulation, which poses substantial difficulties for GUI agents. Conversely, employing a diverse set of APIs for application control offers a more efficient and agent-friendly means of operation. However, this approach introduces its own set of challenges, including the complexity of API implementation, limitations in agent adaptability and flexibility, and the fact that many applications restrict CLI access due to security considerations. To address these issues, we propose an innovative API-GUI paradigm that unifies both action spaces. This design enables agents to leverage the efficiency of API invocation while retaining the versatility of GUI-based operations for general tasks.
Moreover, we develop an automated workflow for desktop application API development. Leveraging LLMs, our framework automates key steps, including requirement analysis, API implementation, and generation of basic test cases. The workflow significantly lowers the barrier for users to create APIs tailored to various applications, thereby facilitating rapid development and deployment.
API Construction In this section, we introduce a methodology for leveraging LLMs to automate API construction. We propose an automated workflow wherein users need only supply exemplar tasks performed within the target application; our LLM then autonomously generates both the necessary API code and corresponding test cases. The workflow comprises three primary stages: requirement analysis, API implementation, and test case generation.
  • Requirement Analysis During the requirements analysis phase, users provide a set of task examples related to the target application as input. The workflow leverages our LLM to analyze these task instances, extracting the essential functionalities required for task completion. It then compares these requirements against the existing API interface definitions to identify potential gaps. If uncovered functionalities are detected, the system automatically generates new API interfaces along with their corresponding parameter specifications.
  • API Implementation Upon obtaining the interface definitions, the workflow systematically iterates over each interface and its associated parameters. For each specification, it leverages the designated Python libraries of the target application to implement the corresponding API functionalities. Additionally, the workflow incorporates error-handling mechanisms and logging to facilitate human debugging and future maintenance. This automated approach not only streamlines API development but also enhances consistency and reusability across different application contexts.
This automated methodology substantially lowers the manual effort required, enabling the creation of application-specific API sets with minimal human intervention. As a result, the barrier for users to develop APIs for diverse applications is significantly reduced. We have developed API sets for multiple commonly used default applications in Ubuntu and integrated them into our Ubuntu virtual machine environment, where their effectiveness has been empirically validated through experiments.
@@ 2.2 Stable Ubuntu Environment for Large-Scale Parallelism
A stable and highly scalable Ubuntu environment is fundamental for constructing behavior cloning datasets and facilitating large-scale RL training. Building on OSWorld (Xie et al., 2024), we have identified several shortcomings of its native implementation. The key challenges identified with the vanilla OSWorld are as follows:
  • Resource Intensiveness and Stability: Virtual machines launched via OSWorld are highly CPU-intensive and consume substantial compute resources. High concurrency on a single node often leads to significant instability, including severe performance degradation and frequent system freezes.
  • Network Bottlenecks: Under heavy workloads, the environment exhibits considerable network communication overhead, resulting in frequent connection failures and even loss of IP addresses. Such network issues critically hinder agent-environment interaction and data logging.
  • Lack of Native Distributed Support: OSWorld does not natively support multi-node clustering, prohibiting efficient distributed deployment and operation across compute clusters.
To address these limitations, we build a robust and massively parallelizable OSWorld infrastructure (see Figure 3), with the following design innovations:
  • Standardized, Decoupled Interface: The environment is refactored via the AgentBench API, providing a unified, modular interface. This abstraction layer enables seamless integration with any learning algorithm or agent architecture, and naturally decouples environment execution from the computational back-end. Consequently, resources can be managed more flexibly, and agents can interact with the environment irrespective of the underlying hardware cluster.
  • Lightweight & Efficient VM Deployment: Our solution utilizes qemu-in-docker  [^1] https://github.com/qemus/qemu to orchestrate Ubuntu VM instances in a containerized manner. We build streamlined VM images that mitigate prior network stack issues, bundle essential utilities directly into the image, and reduce resource requirements through optimizing unneeded services or settings. This significantly reduces per-instance CPU usage, enabling a greater density of concurrent environments.
  • Distributed Multi-Node Clustering: To achieve genuine scalability, we employ a gRPC-based  [^2] https://github.com/grpc inter-process communication protocol that links multiple CPU nodes into a distributed cluster. A unified controller service centrally manages resource allocation, environment orchestration, and inter-node synchronization.
  • Web-based Visualization and Monitoring: Our controller incorporates a versatile web interface, enabling real-time visualization of environment statuses, agent states, and cluster resource allocation. The module dramatically improves usability, operational transparency, and facilitates rapid debugging during large-scale experiments.
Through these technical improvements, our system supports deployment of serveral thousands of concurrent environments on a multi-node CPU cluster, as validated by extensive empirical evaluation. Results confirm our platform’s superior stability, resource efficiency, and scalability, making it an enabling infrastructure for large-scale RL and agent-based research.
@@ 2.3 Full-Asynchronous RL Framework for Efficient Training
In addition to ensuring a stable and scalable environment, an efficient RL training framework is crucial for supporting large-scale model training. Most existing RL frameworks rely on synchronous training paradigms, where rollout collection and parameter updates proceed in alternating phases, resulting in substantial training inefficiencies. To address these limitations, we utilize AgentRL (Zhang et al., 2025) for fully asynchronous RL training with the following core design principles:
  • Dedicated Resource Partitioning: Data collection runs on its resource pool, separated from training. The trainer streams data from the replay engine, so data ingestion and computation do not block each other.
  • Dynamic Batch Sizing: The trainer processes incoming data as soon as enough is available, using flexible batch sizes. This reduces idle time and improves resource efficiency.
  • Modular Component Isolation: Actor, reference, and critic modules run independently from the replay engine, each with dedicated resources. We utilize PyTorch distributed groups and NCCL for efficient parameter sharing, which stabilizes GPU memory usage and enhances throughput.
  • Off-policy Bias Mitigation: We limit the replay buffer size and sync all collected trajectories to the training group after each update. This ensures that each trajectory remains close to the latest policy, thereby minimizing off-policy bias in practice.
Through a stable, high-concurrency desktop environment and the decoupling of training from rollout, we markedly enhance the efficiency of sampling and RL training. Our system achieves a high average power consumption per GPU, reflecting optimal resource utilization. This design supports scalable, high-throughput RL training by enabling dynamic workload balancing, leading to significant improvement in hardware efficiency and overall training throughput.
@@ 3 The ComputerRL Training
Building upon the stable and efficient environment and training framework 2, we establish a robust foundation for large-scale agent training. However, scaling end-to-end training is hindered by challenges in initializing a capable and diverse base policy and by entropy collapse during RL, which limits sustained improvement. In this section, we detail our approach to scalable ComputerRL training, highlighting how these algorithmic innovations facilitate efficient learning in complex desktop environments.
Figure 4: Overview of ComputerRL training. It includes three stages: (1) BC cold-start with trajectories collected via general LLMs; (2) RL with step-level GRPO using verifiable, rule-based rewards; (3) Entropulse, which alternates RL with SFT on successful rollouts to restore entropy and sustain learning.
@@ 3.1 Behavior Cloning Setup
Trajectory Collection with Multiple General LLMs. We manually collect extensive, well-defined tasks and corresponding evaluation functions. However, the large-scale collection of high-quality interaction trajectories remains a fundamental challenge. Manual annotation, while precise, is prohibitively expensive in terms of both human labor and the overhead of developing and maintaining dedicated annotation tools. Furthermore, relying solely on a single model for trajectory generation inevitably results in a limited and homogeneous data distribution, with the attainable quality constrained by the upper bound of that model’s capabilities.
To address these limitations, we leverages the complementary strengths of multiple advanced models to collect a diverse and high-quality set of interaction trajectories. Concretely, our data pipeline consists of three key stages:
@@
    Initial Sampling: For each task curated as described, we utilize closed-source LLMs to sample several trajectories per task independently. We record both the complete interaction trajectories and the outputs produced by the respective evaluation functions. This procedure yields a rich set of diverse trajectories that serve as the foundation for subsequent data augmentation and model adaptation.
@@
    Task-Oriented Fine-Tuning with Stratified Sampling: Following initial data collection, we perform a stratified analysis of task outcomes by categorizing all tasks into three groups based on achieved accuracy:
    • Fully Solved ($\mathrm{acc}=100\%$),
    • Partially Solved ($0\%<\mathrm{acc}<100\%$),
    • Unsolved ($\mathrm{acc}=0\%$).
@@
    Trajectory Augmentation for Hard Tasks: For partially solved tasks, we conduct supervised fine-tuning of our backbone model using the initial trajectories as input. The fine-tuned model is then employed to sample more trajectories for each task, thus substantially expanding the coverage and quality of trajectories for tasks where model proficiency was previously limited.
    For tasks classified as unsolved, we build a model pool consists of high-performing models and randomly pick a model to determine the following action at each step in the interaction process. This approach is motivated by the empirical observation that, while aggregate model performances are comparable, considerable inter-model variance exists at the task level, reflecting distinct areas of expertise. Model collaboration thus enables the generation of trajectories that would be unattainable using any single model in isolation.
We systematically aggregate and filter the collected interaction data, retaining only successful trajectories, and employ them for supervised fine-tuning of the model. This strategy equips the model with robust desktop manipulation capabilities and foundational reasoning abilities, significantly enhancing the performance of the base model.
3.2 Reinforcement Learning from Verifiable Rewards
Step-Level Group Relative Policy Optimization. We have extended the GRPO algorithm (Shao et al., 2024) to the step-level, making it more suitable for agent RL training. For each task $\tau$, the policy $\pi_{\theta}$ interacts with the desktop environment and samples $G$ complete trajectories $\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{G}$. The $i$-th trajectory consists of $L_{i}$ atomic actions (steps) $o_{i,1},\ldots,o_{i,L_{i}}$. All steps from the same task are grouped, and an advantage $A_{i,j}$ is computed for each step. The overall loss aggregates all step advantages as follows:
$\mathcal{J}_{StepGRPO}(\theta)=$ $\;\mathbb{E}_{\mathcal{T}\sim P(\mathcal{T}),\left\{\{o_{i,j}\}_{j=1}^{L_{i}}\right\}_{i=1}^{G}\sim\pi_{\theta_{old}}}\Bigg{[}\frac{1}{\sum_{i=1}^{G}L_{i}}\sum_{i=1}^{G}\sum_{j=1}^{L_{i}}\Bigg{(}\min\Bigg{(}\frac{\pi_{\theta}(o_{i,j}|q)}{\pi_{\theta_{old}}(o_{i,j}|q)}A_{i,j},$
$\qquad\text{clip}\Big{(}\frac{\pi_{\theta}(o_{i,j}|q)}{\pi_{\theta_{old}}(o_{i,j}|q)},1-\epsilon,1+\epsilon\Big{)}A_{i,j}\Bigg{)}-\beta\mathbb{D}_{KL}(\pi_{\theta}\|\pi_{ref})\Bigg{)}\Bigg{]},$

### v2

@@ Abstract
We introduce ComputerRL, a framework for autonomous desktop intelligence that enables agents to operate complex digital workspaces skillfully. ComputerRL features the API-GUI paradigm, which unifies programmatic API calls and direct GUI interaction to address the inherent mismatch between machine agents and human-centric desktop environments. Scaling end-to-end RL training is crucial for improvement and generalization across diverse desktop tasks; however, it remains challenging due to environmental inefficiency and instability during extended training. To support scalable and robust training, we develop a distributed RL infrastructure capable of orchestrating thousands of parallel virtual desktop environments to accelerate large-scale online RL. Furthermore, we propose Entropulse, a training strategy that alternates reinforcement learning with supervised fine-tuning, effectively mitigating entropy collapse during extended training runs. We employ ComputerRL on open models GLM-4-9B-0414 and GLM-4.1V-9B-Thinking, and evaluate them on the OSWorld benchmark. The AutoGLM-OS-9B achieves a new state-of-the-art accuracy of 48.9%, demonstrating significant improvements for general agents in desktop automation. Our code and the new OfficeWorld benchmark are available at our GitHub repository. The algorithm and framework are adopted in building AutoGLM (Liu et al., 2024b).
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023b; Zeng et al., 2022; GLM et al., 2024; Team et al., 2023; Guo et al., 2025; Bai et al., 2023a; Yang et al., 2025a) have dramatically expanded the scope and depth of artificial intelligence capabilities, driving a profound re-examination of our understanding of machine intelligence. Among all scenarios, the emergence of LLM-based GUI (graphical user interface) agents, capable of independently perceiving, reasoning, and executing complex tasks on user devices, has aroused particular interest from researchers (Xi et al., 2023; Wang et al., 2023; Liu et al., 2023). Given that desktops remain central to intelligence-intensive tasks, developing computer use agents is crucial for fundamentally transforming human-computer interactions and elevating AI capabilities (Agashe et al., 2025; Wu et al., 2024a).
@@
Despite previous attempts to develop computer use agents (Agashe et al., 2025; Lei et al., 2024; Xie et al., 2025), enabling them to operate autonomously over extended periods in real-world scenarios remains a significant challenge. The first primary obstacle arises from the fact that GUIs are inherently designed for human interaction, making the simulation of human actions by GUI agents (Liu et al., 2024c; OpenAI, 2025; Qin et al., 2025) a non-trivial and cumbersome endeavor. Second, current mainstream approaches of behavior cloning (BC) (Bain & Sammut, 1995; Bratko et al., 1995), including manual annotation (He et al., 2025) and model distillation (Sun et al., 2024; Xu et al., 2024), are limited in scalability and effectiveness. Manual annotation, while precise, is prohibitively labor-intensive for complex tasks. Model distillation, on the other hand, is constrained by the performance of the teacher models, limiting overall capability. Both methods typically exhibit poor generalization and limited error recovery abilities. Finally, although reinforcement learning (RL) has shown potential for desktop automation tasks (Lu et al., 2025; Feng et al., 2025), its practical application remains restricted due to computational complexity and methodological challenges. Complex environments, slow convergence, and known inefficiencies in RL training (Xie et al., 2024; Bonatti et al., 2024; Yu et al., 2025; Fu et al., 2025) severely limit its large-scale adoption in training computer use agents.
@@
In this work, we propose ComputerRL, an end-to-end algorithmic framework designed to advance desktop-level planning, reasoning, and device operation. This framework includes a new API-GUI interaction paradigm, a scalable RL training infrastructure for computer environments, and an RL algorithm for extended effective training. First, we introduce API-GUI, a large-scale, automatically constructed API ecosystem that enables the agent to transcend the inherent biases of human-oriented operational paradigms. It instead leverages a more machine-oriented approach for device interaction, which combines API calls and GUI actions, thereby significantly enhancing both the versatility and overall performance of the agent. Second, we develop a distributed training infrastructure utilizing virtual machine clusters based on Docker and gRPC protocols for scalability, which is fully compatible with AgentBench (Liu et al., 2023). This infrastructure supports thousands of parallel environments, ensuring high scalability and consistent interactions across all environments. Additionally, we integrate the training infrastructure with the AgentRL framework (Zhang et al., 2025) to facilitate efficient asynchronous training, thereby accelerating the training process. Finally, to counteract stagnation and convergence issues in RL training—specifically, entropy collapse and rising KL divergence—we propose Entropulse, which alternates between RL and SFT phases periodically. This approach maintains exploratory capacity and ensures continuous performance gains (Figure 1(b)).
@@
As a result, by harnessing end-to-end RL and optimization in the desktop environment, ComputerRL has achieved remarkable improvement in understanding and operating GUIs. Evaluation on the OSWorld benchmark (Xie et al., 2024) shows ComputerRL’s significant improvements (see Figure 1(a)) in computer use challenges, achieving a success rate of 48.9% (with 66% performance gain from RL), outperforming other state-of-the-art models including OpenAI CUA o3 (42.9%), UI-TARS-1.5 (42.5%), and Anthropic Claude Sonnet 4 (30.7%).
@@
  • We propose a new interaction paradigm, a shift from human-centric to machine-oriented interaction by introducing a large-scale, automatically constructed API ecosystem integrated with conventional GUI operations. This approach addresses the inherent mismatch between human-designed interfaces and artificial agent capabilities, while achieving superior operational efficiency and generalization performance on computer-based tasks.
@@
  • We introduce Entropulse, a novel training methodology that systematically addresses the challenges of entropy collapse and KL divergence accumulation in extended RL training through strategic alternation between RL and SFT phases, enabling sustained performance improvements and achieving state-of-the-art performance in computer automation.
@@ 2 The ComputerRL Framework
The human-oriented design of GUIs hinders agent efficiency, while limited environment scalability restricts large-scale training. This section presents the ComputerRL framework (see Figure 3), which features an API-GUI paradigm that integrates human-like GUI interactions with efficient API invocation. Additionally, we develop a scalable Ubuntu desktop environment for parallelism and utilize a fully asynchronous RL framework for efficient training.
Figure 3: Overview of ComputerRL framework. We introduce an API-GUI action paradigm that seamlessly integrates automatically constructed APIs with GUI actions to improve agent efficiency and effectiveness. A large-scale parallel desktop environment with 1,000+ real-world instances, combined with an asynchronous RL framework, enables efficient sampling and robust agent training.
@@ 2.1 General API-GUI Paradigm
Existing GUI agents face challenges due to their reliance on human-like interactions, while API-based control offers efficiency but introduces implementation complexity and security restrictions. To address these issues, we propose an API-GUI paradigm that unifies both action spaces, enabling agents to leverage API efficiency while retaining GUI versatility.
We develop an LLM-based automated workflow for application API development (Yang et al., 2024a; Wang et al., 2024), significantly lowering the barrier for API creation. Users provide exemplar tasks, and our system autonomously generates API code and test cases through three stages:
  • Requirement Analysis: Users provide task examples for the target application. Our LLM analyzes these instances, extracts essential functionalities, and compares against existing API interfaces to identify gaps. New interfaces are automatically generated for uncovered functionalities, with a focus on general-purpose functions to minimize complexity and enhance usability.
  • API Implementation: The workflow iterates over each interface definition, implementing API functionalities using designated Python libraries. Error-handling mechanisms and logging are implemented for debugging and maintenance purposes.
  • Test Case Generation: Similar to Li & Yuan (2024), we verify API correctness by checking: (1) runtime error-free invocation and (2) correct results across parameter inputs; failed APIs receive error feedback for autonomous correction.
This methodology enables the creation of application-specific APIs with minimal human intervention. We have developed API sets for multiple Ubuntu applications and validated their effectiveness through experiments. Detailed API development workflow is provided in Appendix A. The agent action space and prompt formulation are detailed in Appendix B and C.
@@ 2.2 Stable Ubuntu Environment for Large-Scale Parallelism
A stable and scalable Ubuntu environment is essential for constructing behavior cloning data and large-scale RL training. Building on OSWorld (Xie et al., 2024), we identify key limitations:
  • Resource Intensiveness and Stability: VMs are CPU-intensive and unstable under high concurrency, causing performance degradation and system freezes.
  • Network Bottlenecks: Heavy workloads cause network overhead, connection failures, and IP address loss, hindering agent interaction and logging.
  • Lack of Native Distributed Support: OSWorld lacks multi-node clustering support, preventing efficient distributed deployment.
To address these limitations, we build a robust and parallelizable OSWorld infrastructure (see Figure 3) with the following innovations:
  • Standardized, Decoupled Interface: We refactor the environment via AgentBench API, providing a unified interface that decouples environment execution from the computational back-end and enables flexible resource management.
  • Lightweight VM Deployment: Using qemu-in-dockere, we deploy containerized Ubuntu VMs with streamlined images that reduce network issues and optimize resource usage, significantly lowering per-instance CPU consumption.
  • Distributed Multi-Node Clustering: We employ gRPC-based communication to link CPU nodes into a distributed cluster with centralized resource allocation and orchestration.
  • Web-based Visualization and Monitoring: A web interface provides real-time visualization of environment statuses, agent states, and resource allocation, improving usability and debugging capabilities.
Through these technical improvements, our system supports deployment of several thousands of concurrent environments on a multi-node CPU cluster, as validated by extensive empirical evaluation. Results confirm our platform’s superior stability, resource efficiency, and scalability, making it an enabling infrastructure for large-scale RL and agent-based research.
@@ 2.3 Full-Asynchronous RL Framework for Efficient Training
Existing RL frameworks rely on synchronous training paradigms, where rollout collection and parameter updates are alternated, resulting in training inefficiencies. To address the limitation, we use the AgentRL framework (Zhang et al., 2025) for fully asynchronous RL training with the following designs:
  • Resource Partitioning: Data collection runs on dedicated resources while the trainer streams data from the replay engine, preventing mutual blocking.
  • Dynamic Batch Sizing: The trainer processes incoming data with flexible batch sizes, reducing idle time and improving efficiency.
  • Modular Component Isolation: Actor, reference, and critic modules run independently with dedicated resources. We utilize PyTorch distributed groups and NCCL for efficient parameter sharing.
  • Off-policy Bias Mitigation: We limit the replay buffer size and sync trajectories after each update, ensuring trajectories remain close to the latest policy.
Through a stable, high-concurrency desktop environment and the decoupling of training from rollout, we markedly enhance the efficiency of sampling and RL training. Our system achieves a high average power consumption per GPU, reflecting optimal resource utilization. This design supports scalable, high-throughput RL training by enabling dynamic workload balancing, resulting in a significant improvement in hardware efficiency and overall training throughput.
@@ 3 The ComputerRL Training
In Section 2, we establish a robust foundation for large-scale agent training. However, scaling end-to-end training still faces challenges in initializing a capable base policy and entropy collapse during RL. This section details our scalable ComputerRL training approach and its algorithmic innovations for extended training in desktop environments.
@@ 3.1 Behavior Cloning Setup
Trajectory Collection with Multiple LLMs. We manually collect extensive tasks with corresponding evaluation functions (see Appendix E) and augment to construct an 8k-task dataset. However, the large-scale collection of high-quality trajectories remains challenging. Manual annotation is prohibitively expensive, and relying on a single model for trajectory generation results in limited and homogeneous data distribution constrained by that model’s capabilities. To address these limitations, we leverage the complementary strengths of multiple advanced models to collect a diverse and high-quality set of interaction trajectories. Concretely, our data pipeline consists of three key stages:
@@
    Initial Sampling: For each task, we utilize closed-source LLMs to sample several trajectories per task independently. We record both the complete interaction trajectories and the outputs produced by the respective evaluation functions. This procedure yields a rich set of diverse trajectories that serve as the foundation for subsequent data augmentation and model adaptation.
@@
    Outcome Stratification: Following initial data collection, we perform a stratified analysis of task outcomes by categorizing all tasks into three groups based on achieved accuracy: Fully Solved ($\mathrm{acc}=100\%$), Partially Solved ($0<\mathrm{acc}<100\%$), and Unsolved ($\mathrm{acc}=0\%$).
@@
    Task-Oriented Augmentation with Stratified Sampling: For partially solved tasks, we conduct SFT on our backbone model using the initial trajectories as input. The fine-tuned model is then used to sample additional trajectories for each task, thereby substantially expanding the coverage and quality of trajectories for tasks where model proficiency was previously limited.
    For tasks classified as unsolved, we build a model pool of high-performing models and randomly select one to determine each action. This approach leverages inter-model variance at the task level, as different models exhibit distinct areas of expertise despite comparable aggregate performance, enabling trajectory generation that is unattainable by any single model.
Figure 4: Overview of ComputerRL, which includes three stages: (1) BC cold start with trajectories collected from general LLMs; (2) RL with step-level GRPO using verifiable, rule-based rewards; (3) Entropulse, which alternates RL with SFT on correct rollouts to restore entropy and sustain learning.
We systematically aggregate and filter the collected interaction data, retaining only successful trajectories (180k+ correct steps), and employ them for supervised fine-tuning of the model. This strategy equips the model with robust desktop manipulation capabilities and foundational reasoning abilities, significantly enhancing the performance of the base model.
3.2 Reinforcement Learning with Verifiable Rewards
Step-Level Group Relative Policy Optimization. We extend the GRPO algorithm (Shao et al., 2024) to the step-level, making it more suitable for agent RL training. For each task $\tau$, the policy $\pi_{\theta}$ interacts with the desktop environment and samples $G$ trajectories $\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{G}$. The $i$-th trajectory consists of $L_{i}$ step-level actions $o_{i,1},\ldots,o_{i,L_{i}}$. All steps from the same task are grouped, and the advantage $A_{i,j}$ is computed for each step. The overall loss aggregates all step advantages as follows:
$\mathcal{J}_{StepGRPO}(\theta)=$ $\;\mathbb{E}_{\mathcal{T}\sim P(\mathcal{T}),\left\{\{o_{i,j}\}_{j=1}^{L_{i}}\right\}_{i=1}^{G}\sim\pi_{\theta_{old}}}\Bigg[\frac{1}{\sum_{i=1}^{G}L_{i}}\sum_{i=1}^{G}\sum_{j=1}^{L_{i}}\Bigg(\min\Bigg(\frac{\pi_{\theta}(o_{i,j}|q_{i,j})}{\pi_{\theta_{old}}(o_{i,j}|q_{i,j})}A_{i,j},$
$\qquad\text{clip}\Big(\frac{\pi_{\theta}(o_{i,j}|q_{i,j})}{\pi_{\theta_{old}}(o_{i,j}|q_{i,j})},1-\epsilon,1+\epsilon\Big)A_{i,j}\Bigg)-\beta\mathbb{D}_{KL}(\pi_{\theta}\|\pi_{ref})\Bigg)\Bigg],$